OpenStudy (kittiwitti1):

If\[3x^{2}-2x+7=0\]then\[\left(x-\frac{1}{3}\right)^{2}=\]My thoughts: 1. solve using quadratic formula 2. input x into the second equation Just checking my work here.

OpenStudy (anonymous):

That would work..can you plug it into the quadratic formula?

OpenStudy (kittiwitti1):

This is the formula yes?\[x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (kittiwitti1):

So -\[x=\frac{-(-2)\pm\sqrt{(-2)^{2}-4(3)(7)}}{2(3)}\]

OpenStudy (kittiwitti1):

Am I on the right track here?

OpenStudy (anonymous):

What do you think? @hartnn

OpenStudy (mathmath333):

try another method which is called squaring.

OpenStudy (kittiwitti1):

squaring?

OpenStudy (anonymous):

Your quadratic formula set-up is correct.

OpenStudy (mathmath333):

yep squaring ,not familiar ?

OpenStudy (kittiwitti1):

I'm only in high school ^^; Thank you @iGreen.

OpenStudy (kittiwitti1):

\[=\frac{2\pm\sqrt{4-84}}{6}=\frac{2\pm\sqrt{-80}}{6}=\frac{2\pm i\sqrt{80}}{6}\]

OpenStudy (kittiwitti1):

Any errors yet? :o

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{3x^2-2x+7=0\hspace{.33em}\\~\\ x^2-2\cdot\dfrac{1}{3}x+\dfrac{7}{3}=0\hspace{.33em}\\~\\ x^2-2\cdot\dfrac{1}{3}x+\left(\dfrac{1}{3}\right)^2-\left(\dfrac{1}{3}\right)^2+\dfrac{7}{3}=0\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (kittiwitti1):

. . . # too complex # teacher never showed us that # no thanks lol

OpenStudy (kittiwitti1):

(but yes, I understand the gist of it)

OpenStudy (mathmath333):

ok !, was easy though

OpenStudy (kittiwitti1):

Well it would take more time on my computer-based exam tomorrow. I'd rather use a formula I'm familiar with -

OpenStudy (kittiwitti1):

(These are NOT exam questions!)

OpenStudy (kittiwitti1):

Have I made any errors so far in my quadratic equation?

OpenStudy (mathmath333):

by the my method was based on this \(\Large (a-b)^2=a^2-2ab+b^2\)

OpenStudy (kittiwitti1):

\[\text{subsize }\sqrt{80}\text{ becomes }4\sqrt{5}\]\[\text{so: }4i\sqrt{5}\]\[\text{which becomes: }\frac{2\pm 4i\sqrt{5}}{6}\]

OpenStudy (kittiwitti1):

I think I misused the word "subsize"

OpenStudy (kittiwitti1):

...and that's as far as I got really

OpenStudy (kittiwitti1):

@iGreen. did I make any errors?

OpenStudy (anonymous):

Looks correct..

OpenStudy (anonymous):

But the problem is, we can't find the square root of a negative number..I'm stumped. @iambatman

OpenStudy (kittiwitti1):

Can I go any further ... I guess not

hartnn (hartnn):

\(\Large x =\frac{2\pm i\sqrt{80}}{6}= \frac{1}{3} \pm i\frac{\sqrt{80} }{6} \\ \Large x -\dfrac{1}{3} = \pm i\frac{\sqrt{80} }{6} \) ok till here ? bdw, mathmath's way was easier, but lets go with whatever you are comfortable with :)

hartnn (hartnn):

and we don't have to find the square root now, we have to find the square, and square of imaginary number is easy :) \(\i^2 =-1\)

OpenStudy (kittiwitti1):

wha- ok-

hartnn (hartnn):

and you just have to square both sides of this : \(\Large x =\frac{2\pm i\sqrt{80}}{6}= \frac{1}{3} \pm i\frac{\sqrt{80} }{6} \\ \Large x -\dfrac{1}{3} = \pm i\frac{\sqrt{80} }{6}\) :)

OpenStudy (kittiwitti1):

@hartnn I was taught that you cannot simplify out half of an addition/subtraction equation

hartnn (hartnn):

whats a "half of addition equation??"

OpenStudy (kittiwitti1):

like \[\frac{8+5}{2}\] becomes\[\frac{4+5}{1}\]

OpenStudy (kittiwitti1):

I meant \[\sqrt{5}\] sorry

hartnn (hartnn):

yes, that is indded incorrect but thats not what i did

OpenStudy (kittiwitti1):

^

hartnn (hartnn):

 ^^ is perfectly valid